# Jizhou Liu Booth School of Business

### v1

@@ Abstract
Randomized experiments are increasingly employed in two-sided markets, such as buyer-seller platforms, to evaluate treatment effects from marketplace interventions. These experiments must reflect the underlying two-sided market structure in their design (e.g., sellers and buyers), making them particularly challenging to analyze. In this paper, we propose a randomization inference framework to analyze outcomes from such two-sided experiments. Our approach is finite-sample valid under sharp null hypotheses for any test statistic and maintains asymptotic validity under weak null hypotheses through studentization. Moreover, we provide heuristic guidance for choosing among multiple valid randomization tests to enhance statistical power, which we demonstrate empirically. Finally, we demonstrate the performance of our methodology through a series of simulation studies.
KEYWORDS: Causal inference; Conditional randomization test; Interference; Two-sided experiments; Multiple Randomization Design.
@@ 1 Introduction
Randomized experiments are increasingly used to assess the impact of policies in various online domains, including social media, online marketplaces, and streaming media platforms (Kohavi et al., 2020; Athey and Luca, 2019; Thomke, 2003). These environments generate complex datasets through two-sided interactions between multiple types of actors, such as viewers and content creators or customers and suppliers. In this context, policy evaluation is complicated due to market interference, where the policy treatment on certain units may affect the outcomes of different units in the environment, including those never directly assigned to treatment (Blake and Coey, 2014; Toulis and Parkes, 2016). Traditional randomized experiments, designed for single-population settings, often fall short in addressing the complexities arising from interference in two-sided markets (Masoero et al., 2024; Johari et al., 2022; Wager and Xu, 2021).
@@
To overcome these challenges, a novel class of experimental designs, termed Multiple Randomization Designs, has been introduced by Bajari et al. (2023) and further developed by Masoero et al. (2024), specifically for marketplace experimentation. For example, in an online marketplace with buyers and sellers, the experimenter might randomly assign half of the buyers and half of the sellers to treatment groups, applying a policy intervention (e.g., free shipping) to transactions where both the buyer and the seller are treated. Transactions where either the buyer or the seller (or both) are untreated would not receive the intervention. This design creates variation in the exposure to treatment across different transactions, which facilitates the analysis of causal effects in the presence of interference.
@@
In this paper, we propose Fisherian-style randomization tests tailored specifically for analyzing two-sided market experiments of this kind (Fisher, 1953). We formulate two sharp null hypotheses on spillover treatment effects that are naturally aligned with the two-sided structure of these experiments. Our approach adopts the conditional randomization testing framework that restricts attention to a set of “focal units” associated with each tested hypothesis (Athey et al., 2018; Basse et al., 2019; Puelz et al., 2021). Within this framework, we propose two valid testing procedures: a permutation-based procedure, which is straightforward to implement but relies on specific symmetry conditions that may not universally hold; and an alternative procedure, valid under arbitrary designs, which requires sampling from a potentially complex conditional randomization space. For a third null hypothesis on total treatment effects, we identify a novel trade-off among multiple valid conditional randomization tests and provide heuristic guidance on designing tests with enhanced statistical power.
Next, we extend our framework to test weak null hypotheses on average spillover and total treatment effects by incorporating studentized statistics into our randomization tests. Studentization ensures the asymptotic validity of our procedures under the weak null hypothesis, while preserving finite-sample validity under the corresponding sharp null, following insights from related literature (Chung and Romano, 2013; DiCiccio and Romano, 2017a; Zhao and Ding, 2021; Wu and Ding, 2021). Our key technical contribution involves extending the proof techniques of Zhao and Ding (2021) and Masoero et al. (2024) to the context of two-sided experiments. Finally, we validate our methodology and theoretical analysis through extensive simulation studies and a synthetic data example. In particular, we empirically examine how the choice of different conditioning events in our randomization procedures impacts statistical power, and show how the empirical relationship agrees with our theory.
@@
The rest of the paper is organized as follows. Section 2 describes the setup and notation. Section 3 presents the main results under sharp null hypotheses. Section 4 discusses the randomization tests for the weak null hypotheses. Section 5 examines the finite-sample behavior of our methods through simulations. Section 6 concludes the paper.
@@ 2 Setup and Notation
We begin by introducing the two-sided randomized design following the two-population buyer-seller example in Masoero et al. (2024). Consider a marketplace comprising $I$ buyers and $J$ sellers, indexed by $i=1,\ldots,I$ and $j=1,\ldots,J$, respectively. In the marketplace, buyers and sellers interact, producing an outcome of interest, such as the total number of transactions between the buyer-seller pair within a given time period. We denote this outcome between buyer $i$ and seller $j$ as $Y_{i,j}\in\mathbb{R}$. We denote the set of buyer-seller pairs as $\mathbb{U}=\{(i,j):i=1,\dots,I,j=1,\dots,J\}$. The $I\times J$ matrix $\mathbf{Y}=[Y_{i,j}]$ captures the outcomes observed in the marketplace. To examine the impact of various interventions, researchers conduct randomized experiments at the level of individual buyer-seller pairs, using a binary treatment assignment, $W_{i,j}\in\{0,1\}$. These interventions might include incentives such as free shipping or discounts on transaction fees. The matrix $\mathbf{W}=[W_{i,j}]\in\{0,1\}^{I\times J}$ denotes the full treatment assignment. We use $Y_{i,j}(\mathbf{w})$ to denote the potential outcome for the pair $(i,j)$ under treatment assignment $\mathbf{w}\in\{0,1\}^{I\times J}$, and $\mathbf{Y}(\mathbf{w})$ for the full $I\times J$ matrix of potential outcomes under assignment $\mathbf{w}$. As usual, the observed outcomes and the potential outcomes are related to treatment assignment by the consistency relationship $\mathbf{Y}=\mathbf{Y}(\mathbf{W})$.
@@
Let $w^{B}=(w_{1}^{B},\dots,w_{I}^{B})^{\top}\in\mathbb{W}^{B}\subseteq\{0,1\}^{I}$ represent a binary treatment assignment for buyers with probability distribution $p^{B}$; and $w^{S}=(w_{1}^{S},\dots,w_{J}^{S})^{\top}\in\mathbb{W}^{S}\subseteq\{0,1\}^{J}$ denote the treatment assignment for sellers with distribution $p^{S}$. An independent two-sided randomized design is then defined by the induced probability distribution of $\mathbf{W}=w^{B}(w^{S})^{\top}\subseteq\{0,1\}^{I\times J}=\mathbb{W}$, as a product of the two marginal distributions, i.e., $p(\mathbf{W})=p^{B}(w^{B})p^{S}(w^{S})$.
@@
In equation (1), we present three examples of treatment assignment matrices $\mathbf{W}_{1}$, $\mathbf{W}_{2}$, and $\mathbf{W}_{3}$, corresponding to buyer-side, seller-side, and two-sided designs, respectively. Treatment assignments for a single population, $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$, can be equivalently represented by the assignment vectors $w^{B}=(0,1,0,0,1)$ and $w^{S}=(1,0,1,0,0)$, as defined in Definition 2.1. The assignment matrix for a two-sided randomized design, $\mathbf{W}_{3}$, is then defined as the element-wise product (Hadamard product) of $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$, $\mathbf{W}_{3}=\mathbf{W}_{1}\odot\mathbf{W}_{2}$, or as the outer product of the two assignment vectors, $\mathbf{W}_{3}=w^{B}(w^{S})^{\top}$ (as specified in Definition 2.1).
$\begin{array}[]{ccc}\mathbf{W}_{1}=\left(\begin{array}[]{ccccc}0&0&0&0&0\\ 1&1&1&1&1\\ 0&0&0&0&0\\ 0&0&0&0&0\\ 1&1&1&1&1\end{array}\right)&\mathbf{W}_{2}=\left(\begin{array}[]{ccccc}1&0&1&0&0\\ 1&0&1&0&0\\ 1&0&1&0&0\\ 1&0&1&0&0\\ 1&0&1&0&0\end{array}\right)&\mathbf{W}_{3}=\left(\begin{array}[]{ccccc}0&0&0&0&0\\ 1&0&1&0&0\\ 0&0&0&0&0\\ 0&0&0&0&0\\ 1&0&1&0&0\end{array}\right)\\ \text{Buyer-side randomization}&\text{Seller-side randomization}&\text{Simple multiple randomization}\end{array}$ (1)
In Example 2.2, we demonstrate the concept of an independent two-sided experiment under complete randomization. In this design, treatment assignments for each population are uniformly distributed, ensuring a fixed fraction of treated units. Another variant of the two-sided design assigns treatments based on independent and identically distributed (i.i.d.) Bernoulli trials for each buyer and seller. As discussed later, both designs enjoy a nice property known as “design symmetry” or “exchangeability”, which motivates and validates the use of permutation tests as a computationally straightforward approach.
@@
For fixed integers, $0<I_{1}<I$ and $0<J_{1}<J$, $I_{0}=I-I_{1}$ and $J_{0}=J-J_{1}$, let $p^{B}\sim\text{Unif}(\mathbb{W}^{B})\text{ and }p^{S}\sim\text{Unif}(\mathbb{W}^{S})$ where $\mathbb{W}^{B}=\{w^{B}\in\{0,1\}^{I}:\sum_{i=1}^{I}w_{i}^{B}=I_{1}\}$ and $\mathbb{W}^{S}=\{w^{S}\in\{0,1\}^{J}:\sum_{i=1}^{J}w_{i}^{S}=J_{1}\}$.
@@
$p(\boldsymbol{W})=\begin{cases}\left(\begin{array}[]{c}I\\ I_{1}\end{array}\right)\left(\begin{array}[]{c}J\\ J_{1}\end{array}\right)&\text{, if }\boldsymbol{W}\in\left\{w^{B}(w^{S})^{\top}:w^{B}\in\mathbb{W}^{B},w^{S}\in\mathbb{W}^{S}\right\}\\ 0&\text{ otherwise. }\end{cases}$
Next, we introduce the concept of local interference, which forms the basis for the two null hypotheses of interest. Under the classical Stable Unit Treatment Value Assumption (SUTVA) with no interference (Rubin, 1972, 1980), the potential outcome can be represented by only the treatment status of a given pair, i.e. $Y_{i,j}(W_{i,j})$. However, recent literature on online marketplace experiments (Blake and Coey, 2014; Toulis and Parkes, 2016; Johari et al., 2022; Masoero et al., 2024) has pointed out that this assumption is untenable as there is likely interference due to a single seller interacting with multiple buyers or a single buyer interacting with multiple sellers. On the other hand, without any restrictions on the potential outcome function, the analysis of potential outcomes would be intractable as the treatment space is exponentially large. To make progress, we follow Masoero et al. (2024) and assume that interference may only occur between the activities of a given buyer or seller in any given buyer-seller pair.
@@
Following Lemma 3.6 in Masoero et al. (2024) and under Assumption 2.1, we can express the potential outcomes in terms of four distinct treatment exposure cases as follows:111 To be more specific, by Assumption 2.1, the potential outcomes can be represented as $Y_{i,j}(W_{i,j},\bar{w}^{B}_{i},\bar{w}^{S}_{j})$, where $\bar{w}^{B}_{i}=\frac{1}{J}\sum_{j=1}^{J}W_{i,j}$ and $\bar{w}^{S}_{j}=\frac{1}{I}\sum_{i=1}^{I}W_{i,j}$ are observed treated fractions. Since $\bar{w}^{B}_{i}=\frac{1}{J}\sum_{j=1}^{J}w_{i}^{B}w_{j}^{S}=w_{i}^{B}p^{S}$ and $\bar{w}^{S}_{j}=w_{i}^{S}p^{B}$, where $p^{S}$ and $p^{S}$ are the treated fractions of sellers and buyers, respectively, we have $Y_{i,j}(W_{i,j},\bar{w}^{B}_{i},\bar{w}^{S}_{j})=Y_{i,j}(w_{i}^{B}w_{j}^{S},w_{i}^{B}p^{S},w_{j}^{S}p^{B})$. Therefore, the potential outcomes are a function of only $w_{i}^{B}$ and $w_{j}^{S}$ whenever our testing procedures do not alter the treated fractions. This justifies the simplified notation of Equation (2).
$Y_{i,j}(\mathbf{W})=Y_{i,j}(w_{i}^{B},w_{j}^{S})=\begin{cases}Y_{i,j}(0,0)&\text{if both buyer and seller are untreated},\\ Y_{i,j}(1,0)&\text{if buyer is treated and seller is untreated},\\ Y_{i,j}(0,1)&\text{if buyer is untreated and seller is treated},\\ Y_{i,j}(1,1)&\text{if both buyer and seller are treated}.\end{cases}$ (2)
@@
$H_{0}^{buyer}:Y_{i,j}(0,0)=Y_{i,j}(1,0),\text{ for all }i=1,\dots,I,j=1,\dots,J.$
@@
$H_{0}^{seller}:Y_{i,j}(0,0)=Y_{i,j}(0,1),\text{ for all }i=1,\dots,I,j=1,\dots,J.$
@@
$H_{0}^{total}:Y_{i,j}(0,0)=Y_{i,j}(1,1),\text{ for all }i=1,\dots,I,j=1,\dots,J.$
@@
Consider a linear additive model for the potential outcomes:
@@
Hypotheses 2.1 to 2.3 can be re-stated as: $H_{0}^{\text{buyer}}:\beta=0$, $H_{0}^{\text{seller}}:\gamma=0$, and $H_{0}^{\text{total}}:\alpha+\beta+\gamma=0$. Here, $\beta$ and $\gamma$ denote the buyer and seller spillover effects, respectively, while $\alpha$ signifies the direct treatment effect.
Finally, we introduce Assumption 2.2, which imposes symmetry on the experimental design. This symmetry allows us to construct permutation tests that are computationally efficient, but it is not necessary to construct valid randomization procedures.
@@
The probability distributions of treatment assignments for buyers, $p^{B}$, and sellers, $p^{S}$, are exchangeable. Specifically, $p^{B}(w_{1}^{B},\dots,w_{I}^{B})=p^{B}(w_{\pi_{B}(1)}^{B},\dots,w_{\pi_{B}(I)}^{B})$ for all permutations $\pi_{B}:[I]\rightarrow[I]$ and $p^{S}(w_{1}^{S},\dots,w_{J}^{S})=p^{S}(w_{\pi_{S}(1)}^{S},\dots,w_{\pi_{S}(J)}^{S})$ for all permutations $\pi_{S}:[J]\rightarrow[J]$.

### v2

# Jizhou Liu PHBS Business School

@@ Abstract
Randomized experiments are increasingly employed in two-sided markets, such as buyer–seller platforms, to evaluate the effects of marketplace interventions. These experiments must reflect the underlying two-sided market structure in their design and can therefore be challenging to analyze. In this paper, we develop a randomization inference framework for outcomes from two-sided experiments, with a focus on testing and inference for two-sided spillover effects. Our approach is finite-sample valid under sharp null hypotheses. Regarding weak null hypotheses, we find that the commonly used Neyman-style studentization does not universally ensure asymptotic validity, and we document how it depends on the specific formulation of the null. We then propose a two-way variance estimator for studentization that restores asymptotic validity. We further propose methods to improve testing power by exploiting the two-sided structure of the problem, which we validate empirically. We demonstrate our methods through a series of simulation studies and an applied example from a network experiment in micro-lending.
KEYWORDS: Causal inference; Conditional Randomization Test; Two-sided Experiments; Multiple Randomization Design.
@@ 1 Introduction
Randomized experiments are increasingly used to assess the impact of policies in various online domains, including online marketplaces, and streaming media platforms (Kohavi et al., 2020; Athey and Luca, 2019; Thomke, 2003). These environments generate complex datasets through two-sided interactions between multiple types of actors, such as viewers and content creators or buyers and sellers. In this context, policy evaluation is complicated due to market interference, where the policy treatment on certain units may affect the outcomes of different units in the market, including those never assigned to treatment (Blake and Coey, 2014; Toulis and Parkes, 2016). Traditional randomized experiments, designed for single-population settings, often fall short in addressing the complexities arising from interference in two-sided markets (Masoero et al., 2026; Johari et al., 2022; Wager and Xu, 2021).
@@
To overcome these challenges, a novel class of experimental designs, termed Multiple Randomization Designs, was developed independently by Masoero et al. (2026); Bajari et al. (2023) and Johari et al. (2022), specifically for marketplace experimentation. For example, in an online marketplace with buyers and sellers, the experimenter might randomly assign half of the buyers and half of the sellers to treatment groups, applying a policy intervention (e.g., free shipping) to transactions where both the buyer and the seller are treated. Transactions where either the buyer or the seller (or both) are untreated would not receive the intervention. This design creates variation in the exposure to treatment across different transactions, which facilitates the analysis of causal effects in the presence of interference.
@@
In this paper, we propose Fisherian-style randomization tests (Fisher, 1953) tailored specifically for analyzing two-sided market experiments of this kind. We formulate sharp null hypotheses on spillover effects that are naturally aligned with the two-sided structure of these experiments. Our approach adopts the conditional randomization testing framework that restricts attention to a set of “focal units” associated with each tested hypothesis (Athey et al., 2018; Basse et al., 2019; Puelz et al., 2021). In this framework, the spillover hypotheses are not sharp in the traditional sense as they cannot be used to impute all potential outcomes. However, with proper conditioning, the hypotheses become sharp and are thus amenable to standard Fisherian-style randomization.
We propose two valid randomization testing procedures: a permutation-based procedure, which is straightforward to implement but relies on specific symmetry conditions on the design that may not universally hold; and an alternative procedure that is valid under arbitrary designs but requires sampling from a potentially complex space. For a third null hypothesis on total treatment effects, we identify a novel validity-power trade-off among multiple candidate tests and provide heuristic guidance on improving power.
Furthermore, we extend our framework to test weak null hypotheses on average spillover and total treatment effects by incorporating studentized statistics into our randomization tests. While studentization is a well-established technique for ensuring asymptotic validity under weak nulls (Chung and Romano, 2013; DiCiccio and Romano, 2017a; Zhao and Ding, 2021; Wu and Ding, 2021; Toulis, 2025), we demonstrate that its effectiveness in two-sided experiments depends crucially on how the null hypothesis is formulated and what variance estimators are used for studentization. Specifically, we show that the commonly used Neyman-style studentization is asymptotically valid under a null hypothesis concerning focal average effects, but it may fail to control size under a global average null due to the additional variance introduced by the two-sided randomization. Then, we propose a modified, two-way variance estimator for studentization that restores asymptotic validity for the global weak null by absorbing the additional sampling variation. Our key technical contribution involves extending the proof techniques of Zhao and Ding (2021) to account for these two-sided dependencies, a finding we validate through extensive simulation studies.
@@
The rest of the paper is organized as follows. Section 2 describes the setup and notation. Section 3 presents the main results under sharp null hypotheses. Section 4 discusses the randomization tests for the weak null hypotheses. Section 5 examines the finite-sample behavior of our methods through simulations. Section 6 illustrates the proposed inference methods in an empirical application based on the experiment conducted in Comola and Prina (2021). Section 7 concludes the paper.
@@ 2 Setup and Notation
We begin by introducing the two-sided randomized design following the two-population buyer-seller example in Masoero et al. (2026). Consider a marketplace comprising $I$ buyers and $J$ sellers, indexed by $i=1,\ldots,I$ and $j=1,\ldots,J$, respectively. In the marketplace, buyers and sellers interact, producing an outcome of interest, such as the total number of transactions between the buyer-seller pair within a given time period. We denote this observed outcome between buyer $i$ and seller $j$ as $Y_{i,j}\in\mathbb{R}$. We denote the set of buyer-seller pairs as $\mathbb{U}=\{(i,j):i=1,\dots,I,j=1,\dots,J\}$. The $I\times J$ matrix $\mathbf{Y}=[Y_{i,j}]$ captures the outcomes observed in the marketplace.
To examine the impact of various interventions, researchers conduct randomized experiments at the level of individual buyer-seller pairs, using a binary treatment assignment, $W_{i,j}\in\{0,1\}$. These interventions might include incentives such as free shipping or discounts on transaction fees. The matrix $\mathbf{W}=[W_{i,j}]\in\{0,1\}^{I\times J}$ denotes the full treatment assignment. We use $Y_{i,j}(\mathbf{w})$ to denote the potential outcome for the pair $(i,j)$ under treatment assignment $\mathbf{w}\in\{0,1\}^{I\times J}$, and $\mathbf{Y}(\mathbf{w})$ for the full $I\times J$ matrix of potential outcomes under assignment $\mathbf{w}$. As usual, the observed outcomes and the potential outcomes are related to treatment assignment by the consistency relationship $\mathbf{Y}=\mathbf{Y}(\mathbf{W})$.
@@
Let $w^{B}=(w_{1}^{B},\dots,w_{I}^{B})^{\top}\in\mathbb{W}^{B}\subseteq\{0,1\}^{I}$ represent a binary treatment assignment for buyers with probability distribution $p^{B}$; and $w^{S}=(w_{1}^{S},\dots,w_{J}^{S})^{\top}\in\mathbb{W}^{S}\subseteq\{0,1\}^{J}$ denote the treatment assignment for sellers with distribution $p^{S}$. An independent two-sided randomized design is then defined by the induced probability distribution of $\mathbf{W}=w^{B}(w^{S})^{\top}\subseteq\{0,1\}^{I\times J}=\mathbb{W}$, where the buyer and seller assignments are independent.
@@
In equation (1), we present three examples of treatment assignment matrices $\mathbf{W}_{1}$, $\mathbf{W}_{2}$, and $\mathbf{W}_{3}$, corresponding to buyer-side, seller-side, and two-sided designs, respectively. Treatment assignments for a single population, $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$, can be equivalently represented by the assignment vectors $w^{B}=(0,1,0,1)$ and $w^{S}=(1,0,1,0,0)$, as defined in Definition 2.1. The assignment matrix for a two-sided randomized design, $\mathbf{W}_{3}$, is then defined as the element-wise product (Hadamard product) of $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$, $\mathbf{W}_{3}=\mathbf{W}_{1}\odot\mathbf{W}_{2}$, or as the outer product of the two assignment vectors, $\mathbf{W}_{3}=w^{B}(w^{S})^{\top}$, as specified in Definition 2.1.
$\begin{array}[]{ccc}\mathbf{W}_{1}=\left(\begin{array}[]{ccccc}0&0&0&0&0\\ 1&1&1&1&1\\ 0&0&0&0&0\\ 1&1&1&1&1\end{array}\right)&\mathbf{W}_{2}=\left(\begin{array}[]{ccccc}1&0&1&0&0\\ 1&0&1&0&0\\ 1&0&1&0&0\\ 1&0&1&0&0\end{array}\right)&\mathbf{W}_{3}=\left(\begin{array}[]{ccccc}0&0&0&0&0\\ 1&0&1&0&0\\ 0&0&0&0&0\\ 1&0&1&0&0\end{array}\right)\\ \text{Buyer-side randomization}&\text{Seller-side randomization}&\text{Two-sided randomization}\end{array}$ (1)
In Example 2.2, we demonstrate the concept of an independent two-sided experiment under complete randomization. In this design, treatment assignments for each population are uniformly distributed, ensuring a fixed fraction of treated units. Another variant of the two-sided design assigns treatments based on independent and identically distributed (i.i.d.) Bernoulli trials for each buyer and seller. As discussed later, both designs enjoy a nice property known as “design symmetry” or “exchangeability”, which motivates the use of permutation tests as a computationally straightforward approach.
@@
For fixed integers, $0<I_{1}<I$ and $0<J_{1}<J$, $I_{0}=I-I_{1}$ and $J_{0}=J-J_{1}$, let $p^{B}\sim\text{Unif}(\mathbb{W}^{B})\text{ and }p^{S}\sim\text{Unif}(\mathbb{W}^{S})$ where $\mathbb{W}^{B}=\{w^{B}\in\{0,1\}^{I}:\sum_{i=1}^{I}w_{i}^{B}=I_{1}\}$ and $\mathbb{W}^{S}=\{w^{S}\in\{0,1\}^{J}:\sum_{j=1}^{J}w_{j}^{S}=J_{1}\}$.
@@
$p(\boldsymbol{W})=\begin{cases}\displaystyle\bigl[C(I,I_{1})\,C(J,J_{1})\bigr]^{-1},&\text{if }\boldsymbol{W}=w^{B}(w^{S})^{\top}\text{ for some }w^{B}\in\mathbb{W}^{B},\;w^{S}\in\mathbb{W}^{S},\\[10.00002pt] 0,&\text{otherwise},\end{cases}$
Next, we introduce the concept of local interference, which forms the basis for the two null hypotheses of interest. Under the classical Stable Unit Treatment Value Assumption (SUTVA) (Rubin, 1972, 1980), the potential outcome can be represented by only the treatment status of a given pair, i.e. $Y_{i,j}(W_{i,j})$. However, recent literature on online marketplace experiments (Blake and Coey, 2014; Toulis and Parkes, 2016; Johari et al., 2022; Masoero et al., 2026) has pointed out that this assumption is untenable as there is likely interference due to a single seller interacting with multiple buyers or a single buyer interacting with multiple sellers. On the other hand, without any restrictions on the potential outcome function, the analysis would be intractable as the treatment space is exponentially large. To make progress, we follow Masoero et al. (2026) and assume that interference may only occur between the activities of a given buyer or seller in any given buyer-seller pair.
@@
Following Lemma 3.5 in Masoero et al. (2026) and under Assumption 2.1, we can express the potential outcomes in terms of four distinct treatment exposure cases as follows: [^1] To be more specific, by Assumption 2.1, the potential outcomes can be represented as $Y_{i,j}(W_{i,j},\bar{w}^{B}_{i},\bar{w}^{S}_{j})$, where $\bar{w}^{B}_{i}=\frac{1}{J}\sum_{j=1}^{J}W_{i,j}$ and $\bar{w}^{S}_{j}=\frac{1}{I}\sum_{i=1}^{I}W_{i,j}$ are observed treated fractions. Since $\bar{w}^{B}_{i}=\frac{1}{J}\sum_{j=1}^{J}w_{i}^{B}w_{j}^{S}=w_{i}^{B}r^{S}$ and $\bar{w}^{S}_{j}=w_{j}^{S}r^{B}$, where $r^{S}$ and $r^{B}$ are the treated fractions of sellers and buyers, respectively, we have $Y_{i,j}(W_{i,j},\bar{w}^{B}_{i},\bar{w}^{S}_{j})=Y_{i,j}(w_{i}^{B}w_{j}^{S},w_{i}^{B}r^{S},w_{j}^{S}r^{B})$. Therefore, the potential outcomes are a function of only $w_{i}^{B}$ and $w_{j}^{S}$ whenever our testing procedures do not alter the treated fractions. This justifies the simplified notation of Equation (2).
[…]
@@
$H_{0}^{\text{buyer}}:Y_{i,j}(0,0)=Y_{i,j}(1,0),\text{ for all }i=1,\dots,I,j=1,\dots,J.$
@@
$H_{0}^{\text{seller}}:Y_{i,j}(0,0)=Y_{i,j}(0,1),\text{ for all }i=1,\dots,I,j=1,\dots,J.$
@@
$H_{0}^{\text{total}}:Y_{i,j}(0,0)=Y_{i,j}(1,1),\text{ for all }i=1,\dots,I,j=1,\dots,J.$
@@
Consider a linear additive model for the potential outcomes (with independent errors):
@@
The null hypotheses in 2.1 to 2.3 can be re-stated as $H_{0}^{\text{buyer}}:\beta=0$, $H_{0}^{\text{seller}}:\gamma=0$, and $H_{0}^{\text{total}}:\alpha+\beta+\gamma=0$. Here, $\beta$ and $\gamma$ denote the buyer and seller spillover effects, respectively, while $\alpha$ is the direct treatment effect.
Finally, we introduce Assumption 2.2, which imposes symmetry on the experimental design. This symmetry allows us to construct permutation tests that are computationally efficient, but it is not necessary to construct valid randomization procedures in general.
@@
The probability distributions of treatment assignments for buyers, $p^{B}$, and sellers, $p^{S}$, are exchangeable. Specifically, $p^{B}(w_{1}^{B},\dots,w_{I}^{B})=p^{B}(w_{\pi_{B}(1)}^{B},\dots,w_{\pi_{B}(I)}^{B})$ for all permutations $\pi_{B}:[I]\rightarrow[I]$ and similarly $p^{S}(w_{1}^{S},\dots,w_{J}^{S})=p^{S}(w_{\pi_{S}(1)}^{S},\dots,w_{\pi_{S}(J)}^{S})$ for all permutations $\pi_{S}:[J]\rightarrow[J]$.